4. Suppose you wish to encrypt the message, M 42 using RSA encryption. Given a public key where p- 23 and q-11 and the relative prime e- 7. Find n, and show all necessary steps to encrypt your message (42). (Hint: check p.411 of the text for information on public key RSA) (5 points)
Answer #1
p = 23, q = 11
M = 42, e = 7

n = p*q = 23*11 = 253

encrypted message = M^e mod n
= (42)^7 mod 253
= 230539333248 mod 253
= 15

Answer: 15